JOINT DIFFERENTIAL GEOMETRY/PDE AND PROBABILITY SEMINAR

Wednesday, March 2, 2005 Padelford C-36 3:50 pm Scattering theory for hyperbolic manifolds

Colin Guillarmou (Purdue University)

we will review results on scattering theory for hyperbolic manifolds and perturbations.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.